find two angles such that the angles are adjacent and form an angle of 120 degress. the larger is 20 degrees less than three times the smaller.
Answer by stanbon(48527)

find two angles such that the angles are adjacent and form an angle of 120 degress. the larger is 20 degrees less than three times the smaller.
-------------------
Equations:
x + y = 120
x = 3y - 20
------------------
Substitute to solve for "Y":
3y -20 + y = 120
4y = 140
y = 35 degrees
------------------
x = 120-y
So, x = 120-35 = 85 degrees
=================================
Cheers,
Stan H.